OpenStudy (anonymous):

Divide 3x2 – 11x – 4 by x – 4.

OpenStudy (anonymous):

divide 3x^2-11x-4 by x-4

OpenStudy (anonymous):

start with \[3x^2-11x+4=(x-4)(3x+1)\]and then cancel

OpenStudy (anonymous):

so i should foil the problem first?

OpenStudy (anonymous):

no no you want to factor first!

OpenStudy (anonymous):

this lesson has nothing to do with factoring, so i dont know if doing that would work

OpenStudy (anonymous):

\[\frac{3x^2 – 11x – 4}{x – 4}=\frac{(x-4)(3x+1)}{x-4}=3x+1\]

OpenStudy (anonymous):

besides dividing, factoring also works?

OpenStudy (anonymous):

yes, in this case you can factor and cancel, but not in every case. sometimes you get a remainder

OpenStudy (anonymous):

can you give me an example of where you cant use this method?